4) What is the equity value of the company given the following

information?

 Current share price: $40.00

 Basic shares outstanding: 400.0mm

 50.0mm options outstanding with an exercise price of $20.00

 5.0mm warrants with an exercise price of $45.00

A. $1,600.0mm

B. $1,500.0mm

C. $1,700.0mm

D. $1,625.0mm
Answer #1

Hi

We will use treasury stock method to calculate diluted shares outstanding.

Here Options are having exercise price of $20, which is lower than current share price, So options are in the money

So total process with options exercise = 50*20 = $1000 Mm

And total shares with $1000 Mm using current share price = 1000/40 = $25 mm

warrants exercise price is more than share price so it is out of the money. Hence it will not be counted in diluted shares

So total shares = 400+25 = $425mm

current share price = $40

so equity value = 425*40 = $1,700 mm

Hence option C is correct answer here.

Thanks